OpenStudy (anonymous):

Negative Exponents: 1. What is the value of the expression 3^-4? a) -81 b) -12 c) -1/81 d) 1/81 2. What is the value of the expression 1/4^3? a) 12 b) 64 c) 1/64 d) 1/12 3. What is the value of the expression 2/4^3? a) 1/32 b) 1/6 c) 128 d) 256 4. What value of K solves the equation, K^-3=1/27? a) -81 b) -9 c) -3 d) 3 5. What value of N solves the equation, 3^N=1/81 a) -243 b) -27 c) -4 d) 3 Thank you so much!!

OpenStudy (john_es):

1) 3^(-4)=1/3^4=1/81 2)1/4^3=1/64 3)2/4^3=2/64=1/32 4) \[k^{-3}=1/27\Rightarrow \frac{1}{k^3}=\frac{1}{3^3}\Rightarrow k=3\] 5) \[3^{N}=1/81\Rightarrow {3^N}=\frac{1}{3^4}=3^{-4}\Rightarrow N=-4\]

OpenStudy (anonymous):

@kerfuffle :Do u need help ???:)

OpenStudy (anonymous):

Look :If we have a negative power have this : 2^-1=1/2^1 9^-3=1/9^3 Got it ??:)